OpenStudy (anonymous):

v^6/v^-2 divide & simplify.

OpenStudy (anonymous):

v^6/v^-2=v^(6-(-2))=v^(6+2)

OpenStudy (anonymous):

v^8

OpenStudy (anonymous):

I see, so you just change the exponents to positive and add them?

OpenStudy (anonymous):

when you divide exponents you subtract, the denominator is a negative exponent. subtracting a negative number is the same as adding the positive number